# Holistic Evaluation of Language Models

### v1

@@ 5 Limitations
Our study has several limitations. First, we focus on widely used frontier LMs rather than open-source models; while this choice highlights that even strong LMs remain sensitive to prompt design, it limits the generality of our findings. Second, our benchmark suite, though diverse, is not exhaustive, and future work should evaluate structured prompting across additional domains. Third, we evaluate a subset of structured prompting methods from the DSPy-family; alternative frameworks could yield higher ceilings. However, our goal is not to identify the optimal prompting method, but to demonstrate that fixed-prompt (without CoT) leaderboard evaluations systematically underestimate LM performance and often distort model comparisons.
@@ 6 Conclusion
By integrating DSPy with HELM, we empirically estimate LM performance ceilings, obtaining more representative estimates. Across seven benchmarks and four LMs, we find that without structured prompting: (i) HELM underestimates LM performance (by 4% average), (ii) performance estimates vary more across benchmarks (++2% standard deviation), (iii) performance gaps are misrepresented (leaderboard rankings flip on $3/7$ benchmarks), and (iv) introducing reasoning (chain-of-thought) reduces LM sensitivity to prompt design (smaller performance $\Delta$ across prompting methods). Sensitivity is heterogeneous: reasoning LMs show marginal gains, whereas some benchmarks for non-reasoning LMs benefit more. Gains are largely agnostic to the structured prompting method. The key driver of improvement is the transition from the baseline prompt to any CoT variant, with Zero-Shot CoT providing the most cost-efficient instantiation. Together, we show that scalable and automated performance ceiling estimation enables more robust, decision-useful benchmarks.
@@
The datasets used in this study are fully open-source, and include: MMLU-Pro (wang2024mmlu), GPQA (rein2024gpqa), GSM8K (cobbe2021gsm8k), MedCalc-Bench (khandekar2024medcalc), Medec (abacha2024medec), HeadQA (vilares2019head), and MedBullets (medbullets). Further distribution is subject to the data sharing agreements stipulated by the original creators.
@@ Acknowledgments
AA is supported by NIH grant R01 HL167974 and ARPA-H contract AY2AX000045. NHS acknowledges support from the Debrah and Mark Leslie endowment for AI in Healthcare, and salary support from Stanford Healthcare. ASC receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H contracts AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Unrelated to this work, ASC receives research support from GE Healthcare, Philips, Microsoft, Amazon, Google, NVIDIA, Stability; has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, and Brain Key; EA receives consulting fees from Fourier Health. This research was, in part, funded by the Advanced Research Projects Agency for Health (ARPA-H). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Government.
@@ Appendix A Bootstrap Few-Shot with Random Search (BFRS)
Random search over few-shot sets.
With module instructions fixed at $I_{i}\!=\!I_{i}^{\mathrm{seed}}$, BFRS draws $R$ independent candidates
@@ Appendix B MIPROv2
where $I_{i}^{(t)}\sim q_{i}(\cdot\mid\mathrm{ctx}_{i})$ are sampled by the proposer LM given context $\mathrm{ctx}_{i}$ (dataset summary, module signature/code, exemplar demos, and tip prompts).

### v2

@@ 5 Limitations
First, we focus on widely used frontier LMs rather than open-source models. While this choice highlights that even strong models remain sensitive to prompt design, it limits the generality of our findings because frontier LMs differ in training data transparency, accessibility, and reproducibility compared to open-source models. Second, our benchmarks primarily involve multiple-choice and short-form reasoning tasks, and results may not generalize to open-ended generation tasks. Third, we evaluate a subset of structured prompting methods from the DSPy family; alternative frameworks could yield higher ceilings. However, our goal is not to identify the optimal prompting method, but to demonstrate that fixed-prompt (without CoT) leaderboard evaluations can systematically underestimate LM performance and often distort model comparisons and rankings.
@@ 6 Conclusion
By integrating DSPy with HELM, we empirically approximate LM performance ceilings, obtaining more representative estimates. Our results show that structured prompting can materially alter benchmark conclusions, shifting relative LM ordering and improving robustness by reducing sensitivity to arbitrary prompt choices. Sensitivity is heterogeneous: reasoning LMs show marginal gains, whereas some benchmarks for non-reasoning LMs benefit more, and gains are largely agnostic to the particular structured prompting method. The key driver of improvement is the transition from the baseline prompt to any CoT variant, with Zero-Shot CoT providing the most cost-efficient instantiation. Future public leaderboards should report performance under multiple structured prompting methods, enabling practitioners to assess achievable performance across prompting styles and make more informed deployment decisions. Together, we show that scalable and automated performance-ceiling approximation enables more robust, decision-useful benchmarks.
@@
The datasets used in this study are fully open-source and include: MMLU-Pro (wang2024mmlu), GPQA (rein2024gpqa), GSM8K (cobbe2021gsm8k), MedCalc-Bench (khandekar2024medcalc), Medec (abacha2024medec), HeadQA (vilares2019head), and MedBullets (medbullets). Further distribution is subject to the data-sharing agreements stipulated by the original creators.
@@ Acknowledgments
AA is supported by NIH grant R01 HL167974 and ARPA-H contract AY2AX000045. NHS acknowledges support from the Debrah and Mark Leslie endowment for AI in Healthcare, and salary support from Stanford Healthcare. ASC receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H contracts AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Unrelated to this work, ASC receives research support from GE Healthcare, Philips, Microsoft, Amazon, Google, NVIDIA, Stability; has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, Brain Key, and Radiology Partners. This research was, in part, funded by the Advanced Research Projects Agency for Health (ARPA-H). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Government.
@@ Appendix A Bootstrap Few-Shot with Random Search (BFRS)
Random search over few-shots.
With instructions fixed at $I_{i}\!=\!I_{i}^{\mathrm{seed}}$, BFRS draws $R$ candidates
@@ Appendix B MIPROv2
where $I_{i}^{(t)}\sim q_{i}(\cdot\mid\mathrm{ctx}_{i})$ are sampled by the proposer LM given context $\mathrm{ctx}_{i}$.